OpenStudy (love_to_love_you):

Which line is parallel to the line that passes through the points (2, -5) and (-4, 1)? y = -x - 5 y= -2/3 x +3 y = 2/3 x - 2 y = x + 5

OpenStudy (harman.singh):

Parallel lines have the same gradients. So if you find the gradient of the given two points, the line parallel to it would have the same gradient.

OpenStudy (love_to_love_you):

How do you find the gradient?

OpenStudy (harman.singh):

You can use this method to find the gradient: 

OpenStudy (love_to_love_you):

m = 1 - (-5) / -4 - 2 = -1

OpenStudy (love_to_love_you):

so y = -x - 5?

OpenStudy (harman.singh):

Correct :-)

OpenStudy (love_to_love_you):

Thank you! I need help with more problems, can i tag you in them?

OpenStudy (harman.singh):

Yea sure.